# Robustly learning the Hamiltonian dynamics of a superconducting quantum processor

### v2

@@ Abstract
The required precision to perform quantum simulations beyond the capabilities of classical computers imposes major experimental and theoretical challenges. The key to solving these issues are highly precise ways of characterizing analog quantum simulators. Here, we robustly estimate the free Hamiltonian parameters of bosonic excitations in a superconducting-qubit analog quantum simulator from measured time-series of single-mode canonical coordinates. We achieve the required levels of precision in estimating the Hamiltonian parameters by maximally exploiting the model structure, making it robust against noise and state-preparation and measurement (SPAM) errors. Importantly, we are also able to obtain tomographic information about those SPAM errors from the same data, crucial for the experimental applicability of Hamiltonian learning in dynamical quantum-quench experiments. Our learning algorithm is highly scalable both in terms of the required amounts of data and post-processing. To achieve this, we develop a new super-resolution technique coined tensorESPRIT for frequency extraction from matrix time-series. The algorithm then combines tensorESPRIT with constrained manifold optimization for the eigenspace reconstruction with pre- and post-processing stages. For up to 14 coupled superconducting qubits on two Sycamore processors, we identify the Hamiltonian parameters—verifying the implementation on one of them up to sub-MHz precision—and construct a spatial implementation error map for a grid of 27 qubits. Our results constitute a fully characterized, highly accurate implementation of an analog dynamical quantum simulation and introduce a diagnostic toolkit for understanding, calibrating, and improving analog quantum processors.
@@
Analog quantum simulators promise to shed light on fundamental questions of physics that have remained elusive to the standard methods of inference Feynman (1982); Lloyd (1996). Recently, enormous progress in controlling individual quantum degrees of freedom has been made towards making this vision a reality Bloch et al. (2012); Blatt and Roos (2012); Smith et al. (2016); Ebadi et al. (2021). While in digital quantum computers small errors can be corrected Aharonov and Ben-Or (2008), it is intrinsically difficult to error-correct analog devices. Yet, the usefulness of analog quantum simulators as computational tools depends on the error of the implemented dynamics. Meeting this requirement hinges on devising characterization methods that not only yield a benchmark of the overall functioning of the device (e.g., Derbyshire et al., 2020; Shaffer et al., 2021; Helsen et al., 2020), but more importantly provide diagnostic information about the sources of errors.
@@
Developing characterization tools for analog quantum simulators requires hardware developments as well as theoretical analysis and method development. With the advent of highly controlled quantum systems, efficient methods for identifying certain Hamiltonian parameters from dynamical data have been devised for specific classes of Hamiltonians. Key ideas are the use of Fourier analysis Schirmer et al. (2004); Cole et al. (2005, 2006a, 2006b); Schirmer et al. (2008); Schirmer and Oi (2009); Oi and Schirmer (2012) and tracking the dynamics of single excitations Burgarth et al. (2009, 2011); Burgarth and Maruyama (2009); Di Franco et al. (2009); Wieśniak and Markiewicz (2010); Burgarth and Yuasa (2012). For general Hamiltonian models, specific algebraic structures of the Hamiltonian terms can be exploited Zhang and Sarovar (2014); Sone and Cappellaro (2017). Generalizing these ideas, a local Hamiltonian can be learned from a single eigenstate or its steady state Garrison and Grover (2018); Qi and Ranard (2019); Chertkov and Clark (2018); Bairey et al. (2019, 2020); Evans et al. (2019) or using quantum-quenches Li et al. (2020a); Czerwinski (2021), an approach dubbed ‘correlation matrix method’ Elben et al. (2023). Alternatively, one can apply general-purpose machine-learning methods Valenti et al. (2019); Bienias et al. (2021); Krastanov et al. (2019); Che et al. (2021); Wilde et al. (2022). More recently, optimal theoretical guarantees have been derived for Hamiltonian learning schemes Yu et al. (2023); Huang et al. (2023); Li et al. (2023) based on Pauli noise tomography Flammia and Wallman (2020); Harper et al. (2020). Crucially, these protocols assume perfect mid-circuit quenches, which—as we find here—can be a limiting assumption in practice.
@@
This recent rapid theoretical development is not quite matched by concomitant experimental efforts. The effectiveness of some of these methods has been demonstrated for the estimation of a small number of coupling parameters of fixed two- and three-qubit Hamiltonians in nuclear magnetic resonance (NMR) experiments Lapasar et al. (2012); Hou et al. (2017); Chen et al. (2021); Zhao et al. (2021). While in NMR, the dominant noise process is decoherence, in tunable quantum simulators such as superconducting qubits, trapped ions or cold atoms in optical lattices, state preparation and measurement (SPAM) errors, as we also demonstrate here, play a central role. Initial steps at characterizing such errors as well as the dissipative Lindblad dynamics for up to two qubits in a superconducting qubit platform have been taken recently Flurin et al. (2020); Samach et al. (2022). Hamiltonian learning of thermal states has recently also been applied in many-body experiments as a means to characterize the entanglement of up to $20$-qubit subsystems whose reduced states are parameterized by the so-called entanglement Hamiltonian Kokail et al. (2021a, b); Joshi et al. (2023). The challenge remains to develop and experimentally demonstrate the feasibility of scalable methods for a robust and precise identification of Hamiltonian dynamics of intermediate-size systems subject to both incoherent noise and systematic SPAM errors.
@@
To this end, building on previous ideas for Hamiltonian identification Burgarth et al. (2011); Zhang and Sarovar (2014), we devise a simple and robust algorithm that exploits the structure of the system at hand. For the identification we make use of quadratically many experimental time-series tracking excitations via expectation values of canonical coordinates. Our structure-enforcing algorithm isolates two core tasks that need to be solved in Hamiltonian identification after suitable pre-processing of the data: frequency extraction and eigenspace reconstruction.
To solve the first task in a robust and structure-specific way, we develop a novel algorithm coined tensorESPRIT, which utilizes ideas from super-resolving, denoised Fourier analysis Roy et al. (1986); Fannjiang (2016); Li et al. (2020b) and tensor networks to extract frequencies from a matrix time-series. For the second task we use constrained manifold optimization over the orthogonal group Abrudan et al. (2009). Crucially, by explicitly exploiting all structure constraints of the identification problem, our method allows us to distinguish and obtain tomographic information about state-preparation and measurement errors. In the quench-based experiment this information renders identification and verification of the dynamics experimentally feasible in the first place. We further support our method development with numerical simulations of different noise effects and benchmark against more direct algorithmic approaches. We find that in contrast to other approaches our method is scalable to larger system sizes out of the reach of our current experimental efforts.
@@
We characterize the Hamiltonian governing analog dynamics of Google Sycamore chips which consist of a two-dimensional array of nearest-neighbour coupled superconducting qubits. Each physical qubit is a non-linear oscillator with bosonic excitations (microwave photons) Carusotto et al. (2020). Using the rotating-wave approximation the dynamics governing the excitations of the qubits in the rotating frame can be well described by the Bose-Hubbard Hamiltonian Yan et al. (2018)
@@
where $a^{\dagger}_{i}$ and $a_{i}$ denote bosonic creation and annihilation operators at site $i$, respectively, $\mu\in\mathbb{R}^{N}$ are the on-site potentials, $J\in\mathbb{R}^{N\times N}$ are the hopping rates between nearest neighbour qubits, and $\eta\in\mathbb{R}^{N}$ are the strength of on-site interactions. The qubit frequency, the nearest-neighbour coupling between them, and the non-linearity (anharmonicity) set $\mu$, $J$, and $\eta$. We are able to tune $\mu$ and $J$ on nanosecond timescales, while $\eta$ is fixed.
@@
The non-interacting part $H(h)$ of the Hamiltonian $H_{\text{BH}}$ can be inferred when initially preparing a state where only a single qubit is excited with a single photon. For initial states with a single excitation, the interaction term vanishes, hence effectively $\eta=0$. Consequently, only the two lowest energy levels of the non-linear oscillators enter the dynamics. Therefore, referring to them as qubits (two-level systems) is precise. Specifically, we identify the parameters $h_{i,j}$ from dynamical data of the following form. We initialize the system in $\left.\left|{\psi_{n}}\right.\right\rangle:=(\mathbbm{1}+a_{n}^{\dagger})\left.\left|{0}\right.\right\rangle^{\otimes N}/{\sqrt{2}}$ and measure the canonical coordinates $x_{m}=(a_{m}+a_{m}^{\dagger})/2$ and $p_{m}=(a_{m}-a_{m}^{\dagger})/(2\mathrm{i})$ for all combinations of $m,n=1,\ldots,N$. In terms of the qubit architecture, this amounts to local Pauli-$X$ and Pauli-$Y$ basis measurements, respectively. We combine the statistical averages over multiple measurements to obtain an empirical estimator for $\langle a_{m}(t)\rangle_{\psi_{n}}=\langle x_{m}(t)\rangle_{\psi_{n}}+\mathrm{i}\,\langle p_{m}(t)\rangle_{\psi_{n}}$. For particle-number preserving dynamics, this data is of the form
@@
We can identify the generator $h$ of the unitary in two steps (Fig. 1(c)), making use of the eigendecomposition of the Hamiltonian (see Methods). In the first step, the time-dependent part of the identification problem is solved, namely, identifying the Hamiltonian eigenvalues (eigenfrequencies). In the second step, given the eigenvalues, the eigenbasis for the Hamiltonian of $h$ is determined. In order to make the identification method noise-robust, we furthermore exploit structural constraints of the model. First, the Hamiltonian has a spectrum such that the time-series data has a time-independent, sparse frequency spectrum with exactly $N$ contributions. Second, the Fourier coefficients of the data have an explicit form as the outer product of the orthogonal eigenvectors of the Hamiltonian. Third, the Hamiltonian parameter matrix is real and has an a priori known sparse support due to the experimental connectivity constraints. These structural constraints are not respected by various sources of incoherent noise, including particle loss and finite shot noise, and coherent noise, in particular the SPAM error. Thereby, an identification protocol that takes these constraints into account is intrinsically robust against various imperfections.
To robustly identify the sparse frequencies from the experimental data, we develop a new super-resolution and denoising algorithm tensorESPRIT that is applicable to matrix-valued time series and uses tensor network techniques in conjunction with super-resolution techniques for scalar data Fannjiang (2016). Achieving high precision in this step is crucial for identifying the eigenvectors in the presence of noise. To robustly identify the eigenbasis, in the second step, we perform least-square optimization of the time-series data under the orthonormality constraint with a gradient descent algorithm on the manifold structure of the orthogonal group Abrudan et al. (2009). Here, we incorporate the connectivity constraint on the coefficient matrix $h$ by making use of regularization techniques Bühlmann and Geer (2011).
@@
The initial and final ramping pulses result in a time-independent, linear transformation at the beginning and end of the time series. It is important to stress that such ramping pulses are expected to be generic in a wide range of experimental implementations of dynamical analog quantum simulations. Robustness of a Hamiltonian identification method against these imperfections is essential for accurate estimates in practice. We can model the effect of such particle number preserving state preparation and measurement (SPAM) errors via linear maps $S$ and $M$, respectively, see the SM for details. This alters our model of the ideal data (3) to
@@ Results.
We implement and characterize different Hamiltonians from time-series data on two distinct quantum Sycamore processors—Sycamore #1 and #2. The Hamiltonians we implement have a fixed overall hopping strength $J_{i,j}=20$ MHz and site-dependent local potentials $\mu_{i}$ on subsets of qubits. Specifically, we choose the local potentials quasi-randomly $\mu_{q}=20\cos(2\pi qb)\,\text{MHz}$, for $q=1,\ldots,N$, where $b$ is a number between zero and one. In one dimension, this choice corresponds to implementing the Harper Hamiltonian, which exhibits characteristic ‘Hofstadter butterfly’ frequency spectra as a function of the dimensionless magnetic flux $b$ Hofstadter (1976).
@@ Conclusion.
Generalizing our two-step approach developed here, we expect a polynomial scaling with the dimension of the diagnosed particle sector and therefore remain efficient for diagnosing two-, three- and four-body interactions, thus allowing to build trust in the correct implementation of interacting Hamiltonian dynamics as a whole. From a broader perspective, with this work, we hope to contribute to the development of a machinery for precisely characterizing and thereby improving analog quantum devices.

### v3

@@ Abstract
The required precision to perform quantum simulations beyond the capabilities of classical computers imposes major experimental and theoretical challenges. The key to solving these issues are precise means of characterizing analog quantum simulators. Here, we robustly estimate the free Hamiltonian parameters of bosonic excitations in a superconducting-qubit analog quantum simulator from measured time-series of single-mode canonical coordinates. We achieve high levels of precision in estimating the Hamiltonian parameters by exploiting a priori knowledge, making it robust against noise and state-preparation and measurement (SPAM) errors. Importantly, we are also able to obtain tomographic information about those SPAM errors from the same data, crucial for the experimental applicability of Hamiltonian learning in dynamical quantum-quench experiments. Our learning algorithm is scalable both in terms of the required amounts of data and post-processing. To achieve this, we develop a new super-resolution technique coined tensorESPRIT for frequency extraction from matrix time-series. The algorithm then combines tensorESPRIT with constrained manifold optimization for the eigenspace reconstruction with pre- and post-processing stages. For up to 14 coupled superconducting qubits on two Sycamore processors, we identify the Hamiltonian parameters—verifying the implementation on one of them up to sub-MHz precision—and construct a spatial implementation error map for a grid of 27 qubits. Our results constitute an accurate implementation of a dynamical quantum simulation that is characterized using a new diagnostic toolkit for understanding, calibrating, and improving analog quantum processors.
@@
Analog quantum simulators promise to shed light on fundamental questions of physics that have remained elusive to the standard methods of inference feynman_simulating_1982; Lloyd. Recently, enormous progress in controlling individual quantum degrees of freedom has been made towards making this vision a reality BlochSimulation; BlattSimulator; MonroeTimeCrystal; ebadi_quantum_2021. While in digital quantum computers small errors can be corrected aharonov_fault-tolerant_2008, it is intrinsically difficult to error-correct analog devices. Yet, the usefulness of analog quantum simulators as computational tools depends on the error of the implemented dynamics. Meeting this requirement hinges on devising characterization methods that not only yield a benchmark of the overall functioning of the device (e.g., derbyshire_randomized_2020; shaffer_practical_2021; helsen_matchgate_2020), but more importantly provide diagnostic information about the sources of errors.
@@
Developing characterization tools for analog quantum simulators requires hardware developments as well as theoretical analysis and method development. With the advent of highly controlled quantum systems, efficient methods for identifying certain Hamiltonian parameters from dynamical data have been devised for specific classes of Hamiltonians. Key ideas are the use of Fourier analysis schirmer_experimental_2004; cole_identifying_2005; cole_identifying_2006; cole_precision_2006; schirmer_physics-based_2008; schirmer_two-qubit_2009; oi_quantum_2012 and tracking the dynamics of single excitations burgarth_coupling_2009; burgarth_indirect_2011; burgarth_indirect_2009; di_franco_hamiltonian_2009; wiesniak_finding_2010; burgarth_quantum_2012. For general Hamiltonian models, specific algebraic structures of the Hamiltonian terms can be exploited zhang_quantum_2014; sone_exact_2017. Generalizing these ideas, a local Hamiltonian can be learned from a single eigenstate or its steady state garrison_does_2018; qi_determining_2019; chertkov_computational_2018; bairey_learning_2019; bairey_learning_2020; evans_scalable_2019 or using quantum-quenches li_hamiltonian_2020; czerwinski_hamiltonian_2021, an approach dubbed ‘correlation matrix method’ elben_randomized_2023. Alternatively, one can apply general-purpose machine-learning methods valenti_hamiltonian_2019; bienias_meta_2021; krastanov_stochastic_2019; che_learning_2021; wilde_scalably_2022. More recently, optimal theoretical guarantees have been derived for Hamiltonian learning schemes yu_robust_2023-1; huang_learning_2023; li_heisenberg-limited_2023 based on Pauli noise tomography flammia_efficient_2020; harper_efficient_2020. Crucially, these protocols assume perfect mid-circuit quenches, which—as we find here—can be a limiting assumption in practice.
@@
This recent rapid theoretical development is not quite matched by concomitant experimental efforts. The effectiveness of some of these methods has been demonstrated for the estimation of a small number of coupling parameters of fixed two- and three-qubit Hamiltonians in nuclear magnetic resonance (NMR) experiments lapasar_estimation_2012; hou_experimental_2017; chen_experimental_2021; zhao_characterizing_2021. While in NMR, the dominant noise process is decoherence, in tunable quantum simulators such as superconducting qubits, trapped ions or cold atoms in optical lattices, state preparation and measurement (SPAM) errors, as we also demonstrate here, play a central role. Initial steps at characterizing such errors as well as the dissipative Lindblad dynamics for up to two qubits in a superconducting qubit platform have been taken recently flurin_using_2020; samach_lindblad_2021. Hamiltonian learning of thermal states has recently also been applied in many-body experiments as a means to characterize the entanglement of up to $20$-qubit subsystems whose reduced states are parameterized by the so-called entanglement Hamiltonian kokail_entanglement_2021; kokail_quantum_2021; joshi_exploring_2023. The challenge remains to develop and experimentally demonstrate the feasibility of scalable methods for a robust and precise identification of Hamiltonian dynamics of intermediate-size systems subject to both incoherent noise and systematic SPAM errors.
@@
To this end, building on previous ideas for Hamiltonian identification burgarth_indirect_2011; zhang_quantum_2014, we devise a simple and robust algorithm that exploits the structure of the system at hand. For the identification we make use of quadratically many experimental time-series tracking excitations via expectation values of canonical coordinates. Our structure-enforcing algorithm isolates two core tasks that need to be solved in Hamiltonian identification after suitable pre-processing of the data: frequency extraction and eigenspace reconstruction.
To solve the first task in a robust and structure-specific way, we develop a novel algorithm coined tensorESPRIT, which utilizes ideas from super-resolving, denoised Fourier analysis RoyPaulrajKailath:1986; Fannjiang:2016:ESPRIT; LiLiaoFannjiang:2019:Super-resolution and tensor networks to extract frequencies from a matrix time-series. For the second task we use constrained manifold optimization over the orthogonal group abrudan_conjugate_2009. Crucially, by explicitly exploiting all structure constraints of the identification problem, our method allows us to distinguish and obtain tomographic information about state-preparation and measurement errors. In the quench-based experiment this information renders identification and verification of the dynamics experimentally feasible in the first place. We further support our method development with numerical simulations of different noise effects and benchmark against more direct algorithmic approaches. We find that in contrast to other approaches our method is scalable to larger system sizes out of the reach of our current experimental efforts.
@@
We characterize the Hamiltonian governing analog dynamics of Google Sycamore chips which consist of a two-dimensional array of nearest-neighbour coupled superconducting qubits. Each physical qubit is a non-linear oscillator with bosonic excitations (microwave photons) PedramReview. Using the rotating-wave approximation the dynamics governing the excitations of the qubits in the rotating frame can be well described by the Bose-Hubbard Hamiltonian yanTunableCouplingScheme2018
@@
where $a^{\dagger}_{i}$ and $a_{i}$ denote bosonic creation and annihilation operators at site $i$, respectively, $\mu\in\mathbb{R}^{N}$ are the on-site potentials, $J\in\mathbb{R}^{N\times N}$ are the hopping rates between nearest neighbour qubits, and $\eta\in\mathbb{R}^{N}$ are the strength of on-site interactions. The qubit frequency, the nearest-neighbour coupling between them, and the non-linearity (anharmonicity) set $\mu$, $J$, and $\eta$. We are able to tune $\mu$ and $J$ on nanosecond timescales, while $\eta$ is fixed for a given setting of $\mu$ and $J$. Hence, the Sycamore chip can be used to implement time evolution under Hamiltonians of the form (1) at various parameter settings and is therefore an analog simulator. In a practical application such as in Ref. roushan_spectroscopic_2017, it is crucial to benchmark how accurately the implemented dynamics is described by a targeted Hamiltonian.
@@
The non-interacting part $H(h)$ of the Hamiltonian $H_{\text{BH}}$ can be inferred when initially preparing a state where only a single qubit is excited with a single photon. For initial states with a single excitation, the interaction term vanishes, hence effectively $\eta=0$. Consequently, only the two lowest energy levels of the non-linear oscillators enter the dynamics. Therefore, referring to them as qubits (two-level systems) is precise. Specifically, we identify the parameters $h_{i,j}$ from dynamical data of the following form. We initialize the system in $\left.\left|{\psi_{n}}\right.\right\rangle\coloneqq(\mathbbm{1}+a_{n}^{\dagger})\left.\left|{0}\right.\right\rangle^{\otimes N}/{\sqrt{2}}$ and measure the canonical coordinates $x_{m}=(a_{m}+a_{m}^{\dagger})/2$ and $p_{m}=(a_{m}-a_{m}^{\dagger})/(2\mathrm{i})$ for all combinations of $m,n=1,\ldots,N$. In terms of the qubit architecture, this amounts to local Pauli-$X$ and Pauli-$Y$ basis measurements, respectively. We combine the statistical averages over multiple measurements to obtain an empirical estimator for $\langle a_{m}(t)\rangle_{\psi_{n}}=\langle x_{m}(t)\rangle_{\psi_{n}}+\mathrm{i}\,\langle p_{m}(t)\rangle_{\psi_{n}}$. For particle-number preserving dynamics, this data is of the form
@@
We can identify the generator $h$ of the unitary in two steps (Fig. 1(c)), making use of the eigendecomposition of the Hamiltonian (see Methods). In the first step, the time-dependent part of the identification problem is solved, namely, identifying the Hamiltonian eigenvalues (eigenfrequencies). In the second step, given the eigenvalues, the eigenbasis for the Hamiltonian of $h$ is determined. In order to make the identification method noise-robust, we furthermore exploit structural constraints of the model. First, the Hamiltonian has a spectrum such that the time-series data has a time-independent, sparse frequency spectrum with exactly $N$ contributions. Second, the Fourier coefficients of the data have an explicit form as the outer product of the orthogonal eigenvectors of the Hamiltonian. Third, the Hamiltonian parameter matrix is real and has an a priori known sparse support due to the experimental connectivity constraints. These structural constraints are not respected by various sources of incoherent noise, including particle loss and finite shot noise, and coherent noise, in particular the SPAM error. Thereby, an identification protocol that takes these constraints into account is intrinsically robust against various imperfections. Importantly, we do not assume that the dynamics of the device is completely governed by a non-interacting, particle-number preserving Hamiltonian of the form (2). We rather impose this as a constraint on the reconstructed Hamiltonian and, as such, identify the best-fit Hamiltonian satisfying the constraint. Our approach thus robustly identifies the non-interacting part of a potentially interacting system.
To robustly identify the sparse frequencies from the experimental data, we develop a new super-resolution and denoising algorithm tensorESPRIT that is applicable to matrix-valued time series and uses tensor network techniques in conjunction with super-resolution techniques for scalar data Fannjiang:2016:ESPRIT. Achieving high precision in this step is crucial for identifying the eigenvectors in the presence of noise. To robustly identify the eigenbasis, in the second step, we perform least-square optimization of the time-series data under the orthonormality constraint with a gradient descent algorithm on the manifold structure of the orthogonal group abrudan_conjugate_2009. Additionally, we can incorporate the connectivity constraint on the coefficient matrix $h$ by making use of regularization techniques Regularization.
@@
The initial and final ramping pulses result in a time-independent, linear transformation at the beginning and end of the time series. It is important to stress that such ramping pulses are expected to be generic in a wide range of experimental implementations of dynamical analog quantum simulations. Robustness of a Hamiltonian identification method against these imperfections is essential for accurate estimates in practice. We can model the effect of such state preparation and measurement (SPAM) errors via linear maps $S$ and $M$, which alters our model of the ideal data (3) to
@@
These linear maps capture the effect of particle-number preserving quenches, as well as the projection of more general channels to the single-particle subspace. Any deviation of the observed experimental dynamics from our model of the data (4) will be visible in the quality of fit.
@@
Predictive power.
If the dynamics of the device is indeed coherent and particle-number preserving, the learned model will allow us to accurately predict the dynamics of the device in the single-particle subspace. If, additionally, interactions are negligible, the predictive power of our model extends to dynamics of more particles. This allows us to benchmark the Sycamore chip as a programmable quantum simulator of the non-interacting part of a Bose-Hubbard model. Accurately predicting the dynamics of many particles requires a generalization of our method to at least the two-particle sector.
@@ Results.
We implement and characterize different Hamiltonians from time-series data on two distinct quantum Sycamore processors—Sycamore #1 and #2. The Hamiltonians we implement have a fixed overall hopping strength $J_{i,j}=20$ MHz and site-dependent local potentials $\mu_{i}$ on subsets of qubits. Specifically, we choose the local potentials quasi-randomly $\mu_{q}=20\cos(2\pi qb)\,\text{MHz}$, for $q=1,\ldots,N$, where $b$ is a number between zero and one. In one dimension, this choice corresponds to implementing the Harper Hamiltonian, which exhibits characteristic ‘Hofstadter butterfly’ frequency spectra as a function of the dimensionless magnetic flux $b$ Hofstadter1976.
@@ Conclusion.
We have demonstrated and custom-tailored our approach here to a superconducting analog quantum simulation platform. It can be applied directly to any bosonic and fermionic analog simulation platform which allows for accurate preparation and measurement of single particle excitations at specific lattice sites. Generalizing our two-step approach developed here, we expect a polynomial scaling with the dimension of the diagnosed particle sector and therefore remain efficient for diagnosing two-, three- and four-body interactions, thus allowing to build trust in the correct implementation of interacting Hamiltonian dynamics as a whole. Furthermore, it is in some cases possible to adapt the method to Hamiltonians with general non-particle number preserving free part. From a broader perspective, with this work, we hope to contribute to the development of a machinery for precisely characterizing and thereby improving analog quantum devices.